OpenStudy (anonymous):

write an equation in slope intercept form for a line that passes through the origin with a slope of 3/2

OpenStudy (sleepyjess):

If the line passes through the origin, what would the y-intercept be?

OpenStudy (anonymous):

0

OpenStudy (alexandervonhumboldt2):

orgin is (0, 0) so the y-intersept is what?

OpenStudy (alexandervonhumboldt2):

correct 0

OpenStudy (sleepyjess):

Correct, are you familiar with the formula for slope-intercept form?

OpenStudy (alexandervonhumboldt2):

so as the slope is 3/2 then lets y=3/2x+?

OpenStudy (alexandervonhumboldt2):

as y-intersept is 0 then y=3/2x+0 y=3/2x